Status Report

ISS On-Orbit Status 16 Jul 2002

By SpaceRef Editor
July 16, 2002
Filed under , ,

All ISS systems continue to function nominally, except as noted previou
sly or below.

Before breakfast, all ISS crewmembers completed another session of the mont
hly Russian medical experiment protocols PZEh MO-7 (calf volume measurement)
and PZEh MO-8 (body mass measurement).  FE-2 Sergei Treschev set up th
e MO-8 equipment and later broke it down and stowed it away.  The prece
ding session, the first for this crew, was on 6/18.

Repair work on the CDRA (carbon dioxide removal assembly) at and behind the
AR (atmosphere revitalization) rack proceeded today, after yesterday’s hard
work by CDR Valery Korzun and FE-1 Peggy Whitson had not gone without a num
ber of unexpected problems.  [Video and narration sent down during t
he R&R (removal and replacement) helped the ground greatly to follow the
activities and provide assistance. Further instructions were uplinked overn
ight based on yesterday’s events.  CDRA procedures were found to requir
e updating based on yesterday’s "surprises".]

Peggy Whitson inspected and activated the Microgravity Science Glovebox
(MSG) prior to supporting the ongoing troubleshooting of the MSG/SUBSA (sol
idification using a baffle in sealed ampoules) complex.  Objective: to
determine why the SAMS II (space acceleration measurement system II) unit is
not communicating via the MSG LAN2 (local area network 2) connection. &nbsp
;Peggy’s intervention on MSG and its MLC laptop was required to enable the g
round to continue failure investigation from earth.  Because of the tro
ubleshooting, today’s originally planned SUBSA activities were deferred to 7

FE-2 Sergei Treschev worked his way through an extensive "hands-on&quo
t; test procedure with the new Nikon D1 electronic still camera, going by up
linked instructions and consulting with a ground specialist on S-band. &nbsp
;[After setting up the equipment yesterday, today he gave the camera syst
ems, its CompactFlash storage card and associated laptop a thorough shakedow
n, first taking pictures inside the Service Module (SM), then photographing
land surfaces on Earth. Selected shot sequences were then stored on the lapt

Later, Treschev cleaned the removable grilles on the gas-liquid heat ex
changers (GZhT) in the FGB, followed by cleaning of FGB ventilation grilles.

CDR Valery Korzun performed a switchover of the Service Module (SM) audio s
ubsystem (STTS) to its primary set.   [The "Voskhod-M&quot
; STTS and its telephone/telegraph audio paths and comm stations were checke
d out by Korzun and Treschev on 6/26, and operated on the backup set for the
last two weeks (since 7/2)..]

Over the past weekend, Whitson worked on the VAJ (vacuum access jumper)
setup and leak test, a task list item.  Ground specialists are current
ly evaluating the pressure readings, in order to better understand the assoc
iated leak characteristics of the VAJ/ISA (international sampling adapter) s

Yesterday’s EXT MDMs (external computers) software upload was successful, a
nd the second half of the s/w upgrade to R2 is done today.  [In its
initial configuration, EXT-1 MDM was powered up and in Primary, running R1.
 First, the R2 files were successfully loaded to the backup MDM (EXT-2)
EEPROM (electrically erasable reprogrammable read-only memory).  EXT-2
was then rebooted from EEPROM, thus running on R2, and subsequently switche
d to Primary.  The procedure is being repeated today for the new backup
MDM.  Final configuration will be EXT-2 MDM powered up and in Primary
running R2, while EXT-1 is powered off, with R2 loaded in EEPROM.  This
is the required initial configuration to support the S1 truss activation pr
ocedure on 9A.]

A huge explosion on the Sun yesterday (7/15) sparked an X-class solar f
lare and hurled a coronal mass ejection into space.  Its source was a r
emarkable sunspot that is currently crossing the face of the Sun. The large
active region stretches 15 Earth-diameters from end-to-end and poses a threa
t for powerful flares, i.e., eruptions of twisted magnetic fields above the
spot, one of which happened yesterday.  In terms of crew dosage, the IS
S residents were not in any danger.

FE-2 Treschev was scheduled for his session of the weekly NTXN Interactions
data collection, filling out a mood, group and journal questionnaire on the
HRF (human research facility) PC.

CDR Korzun performed the weekly inspection of the RED (resistive exercise d
evice), while Sergei Treschev did the weekly TVIS (treadmill with vibration
isolation and stabilization) maintenance.

All crewmembers performed their regular daily physical exercise.

CSA (Canadian Space Agency) reported that the failed Canadarm2 wrist roll (
WR) joint has safely arrived at the contractor’s facility (MDR), and the inv
estigation of its failure mode is underway.

Replacement of the failed FGB battery #6 is scheduled for 7/22.

MCC-Moscow is currently planning to conduct the first of two upcoming Orlan
-EVAs (EVA-7 & EVA-8) from the DC-1 docking compartment in the night hou
rs of 8/15-16.  Use of SSRMS video cameras to cover the spacewalk will
probably not be requested.

The next Soyuz taxi flight, 5S (on 10/22) with a Belgian guest cosmonaut, w
ill carry some experiments that are to be conducted in the U.S. Lab.  T
hus, preparations are to include the POC (Payload Operations Center) in Hunt
sville, and the requisite U.S. training of the crewmember needs to be schedu
led not later than August.

In ten days, the ISS orbit will reach the highest solar beta angle of the y
ear (72 deg).  For power reasons, this will require a pitch angle bias
in XPOP flight attitude. MCC-Moscow was requested to evaluate possible impac
ts on Russian systems and discuss a joint plan.

The current decline of ISS flight altitude was planned to get the station d
own to the optimal altitude for the next Shuttle rendezvous in August. &nbsp
;However, if the launch of Mission 9A is delayed, this projection no longer
applies, and an early and major reboost would become necessary, depending on
9A launch date.  

Today’s targets of the CEO (crew earth observation) program were E. Medi
terranean Dust and Smog
(a slow-moving storm over the western Med was
drawing dust northward off the Libyan coast. On this pass as ISS reached so
uthern Greece, the crew was to try for oblique and limb views to the right o
f track to document this dust event)
, Nile River Delta (althou
gh nadir views were preferred for this target, XPOP may have limited the cre
w to more oblique views. They were to concentrate on the color of the Medite
rranean along the shoreline of the delta, either side of track)
, Ango
lan Biomass Burning
(the crew had an excellent pass for documenting f
ire activity over most of Angola.  Of interest: looking either side of
track to document the frequency and concentration of fire plumes and point s
ources over this region)
, Florida Thunderstorm Anvils (Dynamic
Event Site: Seasonal afternoon thunderstorm activity was expected over the
Florida peninsula this pass. Of interest: using the long lenses of the ESC [
electronic still camera] for oblique and limb views of the thunderstorms to
the left of track, and, where possible, documenting the vertical structure a
nd anvil tops of the cloud formations)
, Lakes of the Eastern Sierra (low light and XPOP attitude were likely to preclude any detailed phot
os of this target. Crew to look to the right of track for context views of t
he sunlit eastern flanks of the southern Sierras and the Mohave Desert. An i
ntense heat wave is exacerbating seasonally dry conditions in this region an
d forest fires should have been visible in the higher elevations)
, Hi
gh Central Andean Glaciers
(cloud conditions were the best they’ve be
en in several weeks here. Of interest: looking to the left of track for cont
ext views to help locate small ice fields and glaciers in the tropical Andes
, Lake Poopo (since this pass clipped the southwestern corner
of this target area, the crew was to look to the left of track for context
views of the Altiplano Basin with its bright salars and a turbid Lake Poopo)

CEO images can be viewed at the website http://eol <>

U.S. and Russian Segment Status (as of 1:37 pm EDT):

Environmental Control and Life Support (ECLSS) and Thermal
Control (TCS):

Elektron O2 generator is powered On (32-amp mode). Vozdukh CO2 s
crubber is ON in MANUAL cycle mode #5 (vacuum pump failed).  U.S. CDRA
CO2 scrubber is Off (valve failed open in one bed).  BMP Harmful Impuri
ties unit: Absorbent bed #1 in Purify mode, bed #2 in Purify mode.

SM Working Compartment:  Pressure (mmHg) — 763, temperature (d
eg C) — 26.7, ppO2 (mmHg) — 157.7, ppCO2 (mmHg) — 3.9.

SM Transfer Compartment:  Pressure (mmHg) — 762, temperature (
deg C) — 20.9.

FGB Cabin:  Pressure (mmHg) — 752, temperature (deg C) — 24.0

Node:  Pressure (mmHg) — 755.34, temperature (deg C) — 24.8 (
shell); ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.

U.S. Lab:  Pressure (mmHg) — 756.”, temperature (deg C) — 24
.8, ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a;

Joint Airlock (Equip. Lock):  Pressure (mmHg) — 756.”, temper
ature (deg C) — 23.8; shell heater temp (deg C) — 23.4, ppO2 (mmHg) — n/a
; ppCO2 (mmHg) — n/a.

PMA-1:  Shell heater temp (deg C) — 27.6

PMA-2:  Shell heater temp (deg C) — 12.8

(Note: Partial pressures ppO2 and ppCO2 in U.S. segment [USOS] not avail
able because MCA [major constituent analyzer] is failed and in Extended Life
[3D a state that preserves mass spectrometer vacuum but produces
no pp data]). MSA (mass spectrometer assembly) and VGA (verification gas as
sembly) were replaced, but some more work needs to be done.

Electrical Power Systems (EPS):

Beta Gimbal Assembly (BGA) 2B in Autotrack mode; BGA 4B in Autotrac
k mode.

SM batteries: Battery #1 reported failed (6/21); all other batteries (7) ar
e in "Partial charge" mode.

FGB: Battery #6 is offline (failed); all other batteries (5) are in 93Partia
l charge” mode.

Plasma Contactor Unit PCU-1 in Standby mode; PCU-2 in Standby mode.

Thermal Control Systems:

Air conditioner SKV-1 is On. SKV-2 is Off.

Command & Data Handling Systems:

C&C-3 MDM is prime, C&C-2 is back-up, and C&C-1 is in s

GNC-1 MDM is prime; GNC-2 is Backup.

LA-1, LA-2 and LA-3 MDMs are all operating.

PL-1 MDM is operational; PL-2 MDM is Off.

APS-1 (automated payload switch #1) and APS-2 are both On.

SM Terminal Computer (TVM): 2 redundant lanes (of 3) operational.

SM Central Computer (TsVM): 3 redundant lanes (of 3) operational.

Attitude Source:

3 CMGs online.

State vector — US GPS (SIGI string 1)

Attitude — Russian segment

Angular rates — US RGA1 (rate gyro assembly 1)

Communications Systems:

All Russian communications & tracking systems are nominal.

S-band is operating nominally.

Ku-band is operating nominally.

Audio subsystem operating nominally.

Video subsystem operating nominally.

MCOR (medium-rate communications outage recorder) is operating nominally.


SSRMS/Canadarm2 at MBS PDGF 1 (mobile base system/power & data
grapple fixture 1) and MBS PDGF 4, with Keep Alive power on both strings (ba
sed on MBS).

MBS:  Operational on both strings.

RWS (robotics workstations): Cupola is Active, Lab RWS is Off.

ISS Orbit (as of this morning, 7:09 am EDT [3D epoch]):

Mean altitude — 390.4 km

Apogee — 396.1 km

Perigee — 384.7 km

Period — 92.4 min.

Inclination (to Equator) — 51.64 deg

Eccentricity — 0.0008454

Orbits per 24-hr. day — 15.59

Altitude decrease — 85 m (mean) in last 24 hours

Solar Beta Angle — 36.2 deg (magnitude increasing)

Revolutions since FGB/Zarya launch (Nov. ’98) — 20859

Current Flight Attitude — XPOP (x-axis perpendicular to orbit plane 3D ”
sun-fixed” [yaw: ~180 deg, pitch: -5.7 deg., roll: 0 deg])

For more on ISS orbit and naked-eye visibility dates/times, see

SpaceRef staff editor.